OpenStudy (anonymous):

1. Which set represents the inequality x 1 in interval notation? (Points : 1) (Infinite, 1] (1, Infinite) [1, Infinite) (Infinite,1)

jimthompson5910 (jim_thompson5910):

the inequality sign is missing

OpenStudy (anonymous):

Oh yes i know that but its because i don't know how to type the inequality sign .... :)

jimthompson5910 (jim_thompson5910):

type ">=" to mean \(\Large \ge\) type "<=" to mean \(\Large \le\)

OpenStudy (paxpolaris):

\[ (-\infty, 1] = x \le1\]\[ (1,\infty) = x >1\]\[ (1,\infty] = x \ge1\]\[ (-\infty, 1) = x <1\]

OpenStudy (paxpolaris):

correction: \[[1,\infty) = x \ge1\]

OpenStudy (anonymous):

ok the question asked which set represents the inequality x >= -1

jimthompson5910 (jim_thompson5910):

So it would be \[\large [-1,\infty)\] basically it's the interval from -1 to infinity. We include -1 with the square bracket [ but we exclude infinity because infinity is not a number and you cannot reach infinity (but you can approach it). So that's why anytime you deal with -infinity or +infinity, you always use a parenthesis.

OpenStudy (anonymous):

Oh okay ! that makes sense thank you !

OpenStudy (anonymous):

so how would i do this one ? :) Which is the solution set of -4 <= 3- y over the whole numbers? {0, 1, 2, 3, 4, 5, 6} {6, 7, 8, 9, 10, …} {7, 8, 9, 10, …} {0, 1, 2, 3, 4, 5, 6, 7}